OpenStudy (anonymous):

z+5 3z-1 4z-4 find the side length of each triangle

OpenStudy (anonymous):

is this a right triangle?

OpenStudy (anonymous):

equalateral 

OpenStudy (anonymous):

z=3 the triangle's sides are 8 units each

OpenStudy (anonymous):

help

OpenStudy (anonymous):

19?

OpenStudy (anonymous):

i need help with 19

OpenStudy (anonymous):

it's an isosceles right triangle with legs of 2x + 6.8 and 8x + 1.4?

OpenStudy (anonymous):

what s the side lemgth of each triangle \

OpenStudy (anonymous):

2x + 6.8 = 8x + 1.4 5.4 = 6x x = .9 s = 2x + 6.8 = 8.6 units